(a) [1[*(2x*y + 4y2) dy dx (b) [ [(y cos(x) + 6) dy dx cos [**(buye* * *) ay ox (a) LiS.*r-* log(4) dy dx (-x log(y)) dy d
Answer #1

a). $ $(22y toga dy de ES [ flexy +449hy}dte 5 ( 124/5dy + ugyay 3d J [()+4(5) Jom -$ (24(70) +41020) die- f + 4) da * °) + $12> $(F-49%Jdres == (1+1) +5 (1+1) = 5) ! }(y cosa + c)dy dz -S (cez giydy tofudayat = (102 (4*)(sin (96) =sin(o) +6(Te-o) -}(1-0) +41 & tillr - 677 si fa com as by die 9.3213 e $ [cageIdy day for fly to day de [altye edeqi @ Subu in o 02. -Ssze (v) da =6Sze*ds - Integrate by parts Segi tg-ffg f= a; g-el frugell @ Szeļdz - ziel ſuelde = relefalby, gal fatty : 9=4 Sencyldy = yeniy-sidy =ylnly) -y 0 equo subo en © 5-2 Hogydy?dy - S-2Cyboys-y)*dx foule dres-99-C1801